# The numerical Amitsur group

### v1

@@ 7.2. Hirzebruch surfaces
Let $X$ be the Hirzebruch surface of degree $e$. In other words, $X$ is a ruled surface over $\mathbb{P}^{1}$ with a section of self-intersection $-e$. Let $H$ be the class of the section and $F$ be the class of a fiber. In this case, $X$ is a toric surface with maximal rays
@@
where $[D_{1}]=F$, $[D_{2}]=H$, $[D_{3}]=F$, and $[D_{4}]=H+eF$. In this case, $\operatorname{Aut}(\Sigma)\cong S_{2}$ interchanges $D_{1}$ and $D_{3}$. Thus $\operatorname{AutP}(X)$ is trivial in this case. We compute that
@@
which has image $\langle 2F,H,H+eF\rangle$ in $\operatorname{Pic}(X)$. We conclude that
$\operatorname{Am}^{T}(X,G)\cong\begin{cases}0&\textrm{if $e$ is odd,}\\ \mathbb{Z}/2\mathbb{Z}&\textrm{if $e$ is even.}\end{cases}$
@@
If $e$ is odd, then $\operatorname{Am}^{\chi}(X,G)$ is trivial. If $e$ is even, then $\operatorname{Am}^{\chi}(X,G)\cong(\mathbb{Z}/2)^{2}$.
@@
are all in $\operatorname{Pic}^{\chi}(X,G)$. If $e$ is odd, then these generate all of $\operatorname{Pic}(X)$ and therefore $\operatorname{Am}^{\chi}(X,G)$ is trivial.
If $e$ is even, then we conclude that $\langle 2H,2F\rangle$ is a subgroup of $\operatorname{Pic}^{\chi}(X,G)$. We will show that, in fact, equality holds. Let $D=aH+bF$ and consider $E:=\chi(D)D=cH+dF$. If $a$ is odd, then $a=2k+1$ and
@@
which is even when $b$ is odd. Thus, if $e$ is even, then $c,d$ are even for all choices of $D$. Thus, $\operatorname{Pic}^{\chi}(X,G)$ never contains $H$, $H+F$ or $F$. ∎
@@ 7.3. The del Pezzo surface of degree 6
Every line bundle in $P$ is contained in $\operatorname{Pic}^{\chi}(X,G)$, but not conversely. Indeed, consider

### v2

@@ 7.2. Hirzebruch surfaces
Let $X$ be the Hirzebruch surface of degree $e$. In other words, $X$ is a ruled surface over $\mathbb{P}^{1}$ with a section of self-intersection e-e. Let $H$ be the class of the section and $F$ be the class of a fiber. In this case, $X$ is a toric surface with maximal rays
@@
where $D_{1}]=F$, $D_{2}]=H$, $D_{3}]=F$, and $D_{4}]=H+eF$. In this case, $\operatorname{Aut}(\Sigma)\cong S_{2}$ interchanges $D_{1}$ and $D_{3}$. Thus $\operatorname{AutP}(X)$ is trivial in this case. We compute that
@@
which has image $\langle 2F,H,H+eF\rangle$ in $\operatorname{Pic}(X)$. We conclude that, for $J=\operatorname{AutP}(X)=1$, we have
$\operatorname{Am}^{T}(X,J)\cong\begin{cases}0&\textrm{if $e$ is odd,}\\ \mathbb{Z}/2\mathbb{Z}&\textrm{if $e$ is even.}\end{cases}$
@@
Let $J=\operatorname{AutP}(X)=1$. If $e$ is odd, then $\operatorname{Am}^{\chi}(X,J)$ is trivial. If $e$ is even, then $\operatorname{Am}^{\chi}(X,J)\cong(\mathbb{Z}/2)^{2}$.
@@
are all in $\operatorname{Pic}^{\chi}(X,J)$. If $e$ is odd, then these generate all of $\operatorname{Pic}(X)$ and therefore $\operatorname{Am}^{\chi}(X,J)$ is trivial.
If $e$ is even, then we conclude that $\langle 2H,2F\rangle$ is a subgroup of $\operatorname{Pic}^{\chi}(X,J)$. We will show that, in fact, equality holds. Let $D=aH+bF$ and consider $E:=\chi(D)D=cH+dF$. If $a$ is odd, then $a=2k+1$ and
@@
which is even when $b$ is odd. Thus, if $e$ is even, then $c,d$ are even for all choices of $D$. Thus, $\operatorname{Pic}^{\chi}(X,J)$ never contains $H$, $H+F$ or $F$. ∎
@@ 7.3. The del Pezzo surface of degree 6
Every line bundle in $P$ is contained in $\operatorname{Pic}^{\chi}(X,G)$, but not conversely. Indeed, consider a group $G$ such that